OpenStudy (anonymous):

Find the limit of the sequence using L'Hôpital's Rule. \[d_n = n^{5}\left(\sqrt[6]{n^{6}+9}-n\right)\]

OpenStudy (anonymous):

\[d_n = n^{5}\left(\sqrt[6]{n^{6}+9}-n\right)\]

OpenStudy (anonymous):

cant i just guess? lol >.<

OpenStudy (anonymous):

lol I wish. maybe im burning out tonight, but I dont even know how to manipulate it into the correct form..

OpenStudy (anonymous):

yeah, me either....i know the answer is probably 0 though. i dont see how to apply L Hopital's Rule

OpenStudy (zarkon):

it's not zero

OpenStudy (anonymous):

It is not 0. fortunately I have unlimited submissions for this assignment.

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} (\sqrt[6]{n^{6}+9}-n)/(n ^{-5})\] This is of the form 0/0 By L'Hopital's: \[\lim_{n \rightarrow \infty} (n^{6}+9)^{-5/6}*6n^5/(-5n ^{-6})\]

OpenStudy (anonymous):

\[\text{Limit}\left[n^5 \left(\sqrt[6]{n^6+9}-n\right),n\to \text{Infinity}\right]=\frac{3}{2} \]Does this help?

OpenStudy (zarkon):

see told you it wasn't 0 ;)

OpenStudy (anonymous):

well, robtobey you are correct but I dont know how you got that.

OpenStudy (anonymous):

wah, i guess i have no guessing sense when it comes to these type of questions, drats >.< lol

OpenStudy (anonymous):

unlike many, I dont come to this site to have others do my homework for me. I wont be here for the final.

OpenStudy (anonymous):

do you see how to get it in the f(x)/g(x) form? if you do the rest is dull arithmeti

OpenStudy (anonymous):

I "cheated". That is a Mathematica 8 Home Edition expression that I posted. WolframAlpha.com might show how they would solve it, step by step.

OpenStudy (anonymous):

aha. nifty.

OpenStudy (anonymous):

In ody's post above, he says its in the form 0/0. i think its actually the form -inf/0, but that is still valid right?

OpenStudy (anonymous):

-inf over 0 is not valid for L'Hopitals (and sorry, in my post i forgot to include the -1)

OpenStudy (anonymous):

-inf over 0 would be (in most cases) just -infinity

OpenStudy (anonymous):

actully I think you are right, it is 0/0. I couldnt read the powers untill i switched to the web app i got this from.

OpenStudy (zarkon):

simplify to this first \[n^{5}\left(\sqrt[6]{n^{6}+9}-n\right)\] \[n^{5}\left(\sqrt[6]{n^{6}\left(1+\frac{9}{n^6}\right)}-n\right)\] \[n^{5}\left(n\sqrt[6]{1+\frac{9}{n^6}}-n\right)\] \[n^{6}\left(\sqrt[6]{1+\frac{9}{n^6}}-1\right)\] \[\frac{\left(\sqrt[6]{1+\frac{9}{n^6}}-1\right)}{n^{-6}}\] use L'Hospitals rule on this

OpenStudy (anonymous):

I could probably take it from there, but that is a lot of work and I already have an answer.